OpenStudy (anonymous):

Evaluate the integral from 5e to 1 of 1/x dx

OpenStudy (mimi_x3):

\[\int \frac{1}{x}dx =ln(x) + C\] Just sub in the limits.

OpenStudy (anonymous):

I mean 1 to 5e

OpenStudy (mimi_x3):

Well, have you atleast tried it?

OpenStudy (anonymous):

but what's 5e

OpenStudy (mimi_x3):

Just sub it in..assuming \(e=2.71...\) fyi; \(e\) is not a variable.

OpenStudy (anonymous):

1+1n5?

OpenStudy (mimi_x3):

\(ln(1) +ln(5e)

OpenStudy (anonymous):

either one?

OpenStudy (mimi_x3):

both

OpenStudy (kirbykirby):

how do you keep your latex code on the same line as regular text?

OpenStudy (mimi_x3):

\[\(\)\]@kirbykirby

OpenStudy (kirbykirby):

ahh coolies :) thanks hehe

OpenStudy (mimi_x3):

:)

OpenStudy (kirbykirby):

For the integral though, you can "pretend" you are integrating from 1 to 5, and you'll get \( ln(5) - ln(1)\) and substitute the "5" with "5e": \( ln(5e)-ln(1)\) but ln(1) = 0 and ln(5e) = ln(5)+ln(e) by log rules and ln(e) = 1 So the final answer is " \( 1+ln(5)\)"

OpenStudy (kirbykirby):

\(e\) is just a number like \(\pi\), you just keep it written that way